I need to find the solution set of the equation x/(x-2) + 5/x(x-2) = 9/2(x-2). I think the LCD is 2x(x-2), and multiplying the numerators by the missing factors of the LCD, I get 2x^2 + 10/LCD = 9x/LCD, but that's as far as I got.
Answer by scott8148(6628)

good job on the LCD ... since this is an equation, you need to treat both sides the same to keep things equal

just multiply the whole equation by the LCD ... 2x^2+10=9x ...or 2x^2-9x+10=0 ... factoring gives (2x-5)(x-2)=0

so x=5/2 and x=2